# Merging Rate of Opinions via Optimal Transport on Random Measures

### v2

@@ 5.2 Normalized generalized gamma process
In this section we study the impact of the discount parameter $\sigma$ of a generalized gamma CRM, that is, the effect of using a normalized generalized gamma CRM instead of a Dirichlet process in model (13). As in the case of Gamma CRMs, we first prove that the posterior is a Cox CRM and find the expression of its Cox scaled CRM (Proposition 16), which is used to express the OT distance between the posteriors in Corollary 17. We then develop an in-depth asymptotic study of the OT distance as the number of observations diverge. Since the Lévy intensities depend on a latent variable $U=U_{n}$, we first need to unravel its asymptotic behaviour, which crucially depends on the number of distinct observations $k=k_{n}$ (Theorem 18). This gives us the means to study its implications on the behaviour on the merging of opinions (Theorem 19 and Proposition 20).
@@
$c=c(\alpha,\sigma,n,k,U)=\alpha(1+U)^{\sigma}+n-k\sigma.$
@@
and $U=U_{n}$ has probability density proportional to $u^{n-1}(1+u)^{k\sigma-n}e^{-\frac{\alpha}{\sigma}(1+u)^{\sigma}}\mathbbm{1}_{(0,+\infty)}(u)$.
@@
with $U=U_{n}$ and $c$ are defined as in Proposition 16.
@@
$\widehat{\mathcal{W}_{\mathrm{AD}}}=\frac{1}{N}\sum_{i=1}^{N}(\mathcal{J}_{1}^{u_{i}}+\mathcal{J}_{2}^{u_{i}}+\mathcal{A}^{u_{i}}),$
@@
To investigate general features of $\mathcal{W}_{\mathrm{AD}}$ as the sample size $n$ grows, one needs to study the asymptotic behaviour of the latent variables $U=U_{n}$. The next result finds the asymptotic distribution of $(1+U)^{\sigma}$ as the number of observations diverge. We focus on $(1+U)^{\sigma}$ instead of $U$ because the former is featured in the expressions of Proposition 17. The core of the proof shows that $(1+U)^{\sigma}$ has a log-concave density, and concentrates around a point of maximum of its log-density. Indeed, with a change of variable one finds that the density of $(1+U)^{\sigma}$ is proportional to $\exp(-f_{n}(x))$, where
@@
is a convex function. If $f_{n}$ were proportional to $n$, say $f_{n}(x)=ng(x)$ for some smooth function $g$, then the result would come directly from Laplace’s method. Here the situation is more delicate as there are different scales (namely, the constant $\alpha/\sigma$, $k_{n}$ and $n$) in the function $f_{n}$. We extend Laplace’s method to this setting in the supplementary material, which yields a general result of independent interest. Note on the other hand that the density of $U$ itself is not log-concave.
@@
  • $a_{n}\ll b_{n}$ if $(a_{n}/b_{n})_{n\in\mathbb{N}}$ converges to 00. Equivalently, $a_{n}=o(b_{n})$.
@@
$r_{n}=\begin{cases}\alpha^{-\frac{\sigma}{1+\sigma}}\ n^{\frac{\sigma}{1+\sigma}}&\text{\qquad if }k\ll n^{\frac{\sigma}{1+\sigma}},\\ \gamma\ n^{\frac{\sigma}{1+\sigma}}&\text{\qquad if }k\sim\lambda n^{\frac{\sigma}{1+\sigma}},\\ \sigma/\alpha\ k&\text{\qquad if }k\gg n^{\frac{\sigma}{1+\sigma}},\end{cases}$
@@
Theorem 18 highlights a transition in the asymptotic behaviour of $(1+U)^{\sigma}$ as the number of unique values $k$ increases. Indeed, one can easily derive that
@@
This implies that the growth rate of $(1+U)^{\sigma}$ is not affected by $k$ if $k$ does not diverge or diverge sufficiently slow, otherwise it coincides with $k$ itself, as shown in Figure 3.
We now use Theorem 18 to study the asymptotic behaviour of the distance in Proposition 17.
@@
Figure 4: Left: Unique values $k=k_{n}$ for a conditionally i.i.d. sample from a Dirichlet process of parameters $(\bar{\alpha}=1,\bar{P}_{0}=\mathcal{N}(0,1))$, as the number of observations $n$ increases. Right: $\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2*}_{\mathcal{S}}))$ when $\tilde{\mu}^{1}$ is a generalized gamma$(\alpha_{1}=100,P_{0}^{1}=\mathcal{N}(0,1),\sigma)$ CRM and $\tilde{\mu}^{2}$ is a gamma$(\alpha_{2}=100,P_{0}^{2}=\mathcal{N}(0,1))$ CRM, as the number of observations $n$ increases exponentially. The observations are conditionally i.i.d. sample from a Dirichlet process of parameters $(\bar{\alpha}=1,\bar{P}_{0}=\mathcal{N}(0,1))$.
Figure 5: Left: Unique values $k=k_{n}$ for a conditionally i.i.d. sample from a Pitman-Yor process of parameters $(\bar{\alpha}=1,\bar{\sigma}=0.9,\bar{P}_{0}=\mathcal{N}(0,1))$, as the number of observations $n$ increases. Right: $\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2*}_{\mathcal{S}}))$ when $\tilde{\mu}^{1}$ is a generalized gamma$(\alpha_{1}=100,P_{0}^{1}=\mathcal{N}(0,1),\sigma)$ CRM and $\tilde{\mu}^{2}$ is a gamma$(\alpha_{2}=100,P_{0}^{2}=\mathcal{N}(0,1))$ CRM, as the number of observations $n$ increases exponentially. The observations are conditionally i.i.d. sample from a Pitman-Yor process of parameters $(\bar{\alpha}=1,\bar{\sigma}=0.9,\bar{P}_{0}=\mathcal{N}(0,1))$.
@@
Let $\tilde{\mu}_{1},\tilde{\mu}_{2}$ as in Proposition 17 and let $(x_{n})_{n\geq 1}$ be s.t. $k_{n}\ll n$. Then as $n\rightarrow+\infty$,
@@
From a statistical perspective, Theorem 19 detects the regimes where using a normalized generalized gamma CRM will actually guarantee a different learning outcome, with the following take-home message: either the number of observations $n$ is sufficiently small or the number of distinct values $k$ should be sufficiently large, and how large it must be depends on the choice of discount $\sigma$, namely, $k>n^{\frac{\sigma}{1+\sigma}}$. As notable examples, this happens when the observations are i.i.d. from a continuous distribution or when they are conditionally i.i.d. given a normalized generalized gamma CRM or a Pitman-Yor process, but it does not happen when the observations are i.i.d. from a discrete distribution with a finite number of atoms or when they are conditionally i.i.d. given a Dirichlet process or a hierarchical Dirichlet process.
@@ Acknowledgments
The authors would like to thank A. Lijoi and I. Prünster for valuable insights. Part of this research has been done during a visit of the second author to the Department of Statistics at the University of Warwick, which is warmly acknowledged for the hospitality. The second author is also affiliated to the Bocconi Institute for Data Science and Analytics (BIDSA). M. Catalano was partially supported by a Research Development Fund from the University of Warwick.
Supplementary material
Contains the proofs of the results of the present work.

### v3

@@ 5.2 Normalized generalized gamma process
In this section we study the impact of the discount parameter $\sigma$ of a generalized gamma CRM, that is, the effect of using a normalized generalized gamma CRM instead of a Dirichlet process in model (13). As in the case of Gamma CRMs, we first prove that the posterior is a Cox CRM and find the expression of its Cox scaled CRM (Proposition 16), which is used to express the OT distance between the posteriors in Corollary 17. We then develop an in-depth asymptotic study of the OT distance as the number of observations diverge. Since the Lévy intensities depend on a latent variable $U=U_{n}$, we first need to unravel its asymptotic behavior, which crucially depends on the number of distinct observations $k=k_{n}$ (Theorem 18). This gives us the means to study its implications on the behavior on the merging of opinions (Theorem 19 and Proposition 20).
@@
$C=C(\alpha,\sigma,n,k,U)=\alpha(1+U)^{\sigma}+n-k\sigma,$ (20)
where we recall that $U=U_{n}$ has p.d.f. (18).
@@
and $U=U_{n}$ has probability density proportional to $u^{n-1}(1+u)^{k\sigma-n}e^{-\frac{\alpha}{\sigma}(1+u)^{\sigma}}\mathbbm{1}_{(0,+\infty)}(u)$ while $C=C_{n}$ is defined in (20).
@@
with $U=U_{n}$ and $C=C_{n}$ are defined as in Proposition 16.
@@
$\hat{\mathcal{W}}_{\mathrm{AD}}=\frac{1}{N}\sum_{i=1}^{N}(\mathcal{J}_{1}^{u_{i}}+\mathcal{J}_{2}^{u_{i}}+\mathcal{A}^{u_{i}}),$
@@
To investigate general features of $\mathcal{W}_{\mathrm{AD}}$ as the sample size $n$ grows, one needs to study the asymptotic behavior of the latent variables $U=U_{n}$. The next result finds the asymptotic distribution of $(1+U)^{\sigma}$ as the number of observations diverge. We focus on $(1+U)^{\sigma}$ instead of $U$ because the former is featured in the expressions of Proposition 17. The core of the proof shows that $(1+U)^{\sigma}$ has a log-concave density, and concentrates around a point of maximum of its log-density. Indeed, with a change of variable one finds that the density of $(1+U)^{\sigma}$ is proportional to $\exp(-f_{n}(x))$, where
@@
is a convex function. If $f_{n}$ were proportional to $n$, say $f_{n}(x)=ng(x)$ for some smooth function $g$, then the result would come directly from Laplace’s method. Here the situation is more delicate as there are different scales (namely, the constant $\alpha/\sigma$, $k_{n}$ and $n$) in the function $f_{n}$. We extend Laplace’s method to this setting in the Supplementary Material (Catalano and Lavenant, 2025), which yields a general result of independent interest. Note on the other hand that the density of $U$ itself is not log-concave.
@@
  • $a_{n}\ll b_{n}$ if $(a_{n}/b_{n})_{n\in\mathbb{N}}$ converges to 0. Equivalently, $a_{n}=o(b_{n})$.
@@
$r_{n}=\begin{cases}\alpha^{-\frac{\sigma}{1+\sigma}}\ n^{\frac{\sigma}{1+\sigma}}&\text{\qquad if }k\ll n^{\frac{\sigma}{1+\sigma}},\\ \gamma\ n^{\frac{\sigma}{1+\sigma}}&\text{\qquad if }k\sim\lambda n^{\frac{\sigma}{1+\sigma}},\\ \sigma\alpha^{-1}\ k&\text{\qquad if }k\gg n^{\frac{\sigma}{1+\sigma}},\end{cases}$
@@
Theorem 18 highlights a transition in the asymptotic behavior of $(1+U)^{\sigma}$ as the number of unique values $k$ increases. Indeed, one can easily derive that
@@
This implies that the growth rate of $(1+U)^{\sigma}$ is not affected by $k$ if $k$ does not diverge or diverge sufficiently slowly; otherwise it coincides with $k$ itself, as shown in Figure 3.
We now use Theorem 18 to study the asymptotic behavior of the distance in Proposition 17.
@@
[…]
@@
Let $\tilde{\mu}_{1},\tilde{\mu}_{2}$ as in Proposition 17 and let $(x_{n})_{n\geq 1}$ be such that $k_{n}\ll n$. Then as $n\rightarrow+\infty$,
@@
From a statistical perspective, Theorem 19 detects the regimes where using a normalized generalized gamma CRM will actually guarantee a different learning outcome, with the following take-home message: either the number of observations $n$ is sufficiently small or the number of distinct values $k$ should be sufficiently large, and how large it must be depends on the choice of discount $\sigma$, namely, $k\gtrsim n^{\frac{\sigma}{1+\sigma}}$. As notable examples, this happens when the observations are i.i.d. from a continuous distribution or when they are conditionally i.i.d. given a normalized generalized gamma CRM or a Pitman-Yor process, but it does not happen when the observations are i.i.d. from a discrete distribution with a finite number of atoms or when they are conditionally i.i.d. given a Dirichlet process or a hierarchical Dirichlet process.
Remark 6.
Gamma and generalized gamma random measures are arguably the most used classes of completely random measures but specific needs could lead to the use of other CRMs not belonging to this class. In such a case, the merging could be explored with techniques similar to those exposed in this section. Whenever the Lévy measure is available in analytical form, we expect an analogous of Proposition 16 to be readily available. The difficult part is to find the asymptotic behavior of the latent variable $U$. Our analysis is based on an ad hoc strategy that recognizes a log-concave distribution for a transformation of $U$ that appears in our expressions. We currently do not know to what extent this feature can be extended to the latent variables arising from other CRMs.
@@ Acknowledgments
The authors would like to thank the Editor, an Associate Editor, and three Referees for the constructive feedback, which has led to significant improvement to the present work. The authors would also like to thank A. Lijoi and I. Prünster for their valuable insights. Part of this research was conducted at the Department of Statistics at the University of Warwick, during a visit by H. Lavenant and while M. Catalano was a Harrison Early Career Assistant Professor. The Department is warmly acknowledged for its hospitality. H. Lavenant is also affiliated with the Bocconi Institute for Data Science and Analytics (BIDSA). M. Catalano was partially supported by a Research Development Fund from the University of Warwick.